# Boosting Memory Efficiency in Transfer Learning for High-Resolution Medical Image Classification

### v2

@@ Abstract
The success of large-scale pre-trained models has established fine-tuning as a standard method for achieving significant improvements in downstream tasks. However, fine-tuning the entire parameter set of a pre-trained model is costly. Parameter-efficient transfer learning (PETL) has recently emerged as a cost-effective alternative for adapting pre-trained models to downstream tasks. Despite its advantages, the increasing model size and input resolution present challenges for PETL, as the training memory consumption is not reduced as effectively as the parameter usage. In this paper, we introduce Fine-grained Prompt Tuning plus (FPT+), a PETL method designed for high-resolution medical image classification, which significantly reduces the training memory consumption compared to other PETL methods. FPT+ performs transfer learning by training a lightweight side network and accessing pre-trained knowledge from a large pre-trained model (LPM) through fine-grained prompts and fusion modules. Specifically, we freeze the LPM of interest and construct a learnable lightweight side network. The frozen LPM processes high-resolution images to extract fine-grained features, while the side network employs corresponding down-sampled low-resolution images to minimize the memory usage. To enable the side network to leverage pre-trained knowledge, we propose fine-grained prompts and fusion modules, which collaborate to summarize information through the LPM’s intermediate activations. We evaluate FPT+ on eight medical image datasets of varying sizes, modalities, and complexities. Experimental results demonstrate that FPT+ outperforms other PETL methods, using only 1.03% of the learnable parameters and 3.18% of the memory required for fine-tuning an entire ViT-B model. Our code is available online.
@@
Parameter-efficient transfer learning, Memory-effificent transfer learning, Large-scale pre-trained models, High-resolution medical image classification.
@@ I Introduction
In natural images, the objects of interest typically occupy a significant portion of the image and exhibit clear and noticeable attributes (e.g., shape and color). In contrast, the diagnostic cues within medical images are often tiny and spread throughout the entire visual field, making fine-grained information crucial for accurate medical image analysis [25]. Providing such fine-grained information often requires high-resolution input images [24]. However, as illustrated in Fig. 2, this preference for high resolution comes at the cost of increased GPU memory consumption and higher training expenses. For example, a ViT with a patch size of 16 divides a 224 $\times$ 224 image into 196 patches, whereas a 512 $\times$ 512 image generates 1024 patches. This sharp rise in computational demand makes fine-tuning LPMs both expensive and, in some cases, infeasible on devices with limited GPU memory.
@@
Additionally, considering that long input sequences from high-resolution images result in substantial intermediate activations, which can be computationally expensive, FPT+ employs important tokens selection to mitigate this issue. Based on the prior knowledge that medical images of the same modality often share similar anatomical characteristics and have small objects of interest (e.g., lesions and vessels in fundus images), FPT+ selectively fuses features from the intermediate activations to reduce GPU memory consumption. Preloading techniques are also proposed to improve the training efficiency by pre-storing the LPM’s intermediate features.
@@
  • We evaluate different LPMs for transfer learning, including both supervised and self-supervised methods, emphasizing the generalizability of FPT+ across different LPM scales and different pre-trained weights.
This paper extends our previous work FPT [23]. We systematically investigate the impact of several key components in FPT and refine the entire framework to make it simpler and more efficient. Specifically, our extension is presented in several aspects: (1) We reduce the number of layers in the side network and its corresponding fusion modules to significantly decrease memory consumption; (2) We redesign the fusion module by replacing in-projector with layer normalization and align the initial hidden dimension of the fine-grained prompts with the frozen LPM; (3) All fusion modules share the same weights for the out-projector, further reducing the number of learnable parameters; (4) The new design improves the effectiveness of learning pre-trained knowledge, allowing smaller low-resolution inputs for the side network, thus further reducing memory consumption; (5) To verify the generalizability of FPT+, we validate the framework using four additional datasets; (6) We evaluate the scalability and feasibility of FPT+ across pre-trained models of different scales and different weights; (7) Additional efficiency metrics, including training time, throughput, and power usage, are evaluated to comprehensively assess FPT+’s resource consumption and training efficiency; (8) Extensive ablation studies are conducted to qualitatively and quantitatively analyze the impact of each component in FPT+ on performance, parameter usage and memory efficiency.
@@ II-A Parameter-efficient Transfer Learning
Unlike other PETL methods, FPT [23] and our proposed FPT+ aim to improve both parameter and memory efficiency during transfer learning, which is a specific and common requirement in the field of medical image analysis.
@@ II-B PETL for Medical Image Analysis
To the best of our knowledge, our previous work, FPT [23], is the first PETL method specifically designed for high-resolution input settings, simultaneously enhancing both parameter and memory efficiency. We extend FPT by simplifying the entire framework to make it more efficient and more effective. Specifically, we halve the number of layers in the side network and the corresponding fusion modules, and redesign the fusion module’s in-projector. Unlike other PETL methods for medical image analysis, our proposed FPT+ framework aims to simultaneously address the challenge of unaffordable parameter as well as memory utilization when transferring LPMs to high-resolution medical image classification tasks.
@@ III-C Asymmetric Input
Fine-grained information holds significant importance in medical image analysis. To extract such detailed information, one straightforward approach involves increasing the resolution of the input image. However, due to the mechanics of input patchification and the self-attention layer in the ViT architecture, increasing the input resolution leads to a dramatic rise in memory consumption and training cost. For instance, in a ViT with a patch size of 16, increasing the input resolution from 224 to 512 results in an expansion of the input sequence length from 196 to 1024. Furthermore, the dimensions of the self-attention maps in the transformer block escalate from $196^{2}$ to $1024^{2}$. This considerable training overhead presents challenges for fine-tuning on high-resolution datasets.
Although high-resolution inputs make the training cost unfeasible, employing a pre-trained model for inference with high-resolution images remains practical. We here propose an asymmetric input strategy within the FPT+ framework. Specifically, the input images $I_{h}$ for the frozen pre-trained model are of high resolution, whereas the images fed to the learnable side network are low-resolution images $I_{l}$ that get down-sampled from the high-resolution ones. Therefore, the Equation 5 becomes
@@ III-D Fine-grained Prompts and Fusion Modules
where notations for the input source and the layer index are omitted for simplification. The $\text{CA}(\cdot)$ and $\text{LN}(\cdot)$ denote the cross-attention module and the layer normalization operation, respectively. The [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the concatenation operation, and the out-projector $f_{out}$ is a linear layer to align the hidden dimension of the fine-grained prompts to that of the side network. The out-projector is shared across all fusion modules to further increase efficiency. Notably, the number of the fine-grained prompts is considerably smaller than that of the input sequence, leading to reduced training cost. They are removed after each layer’s forward processing in the side network to maintain consistency of the sequence length across layers.
@@ III-F Fine-grained Features Preloading
Supervised (ImageNet)
Full fine-tuning 100 23,128 89.79 ± 0.76 86.87 ± 0.53 96.65 ± 0.29 83.22 ± 1.02 92.49 ± 0.34 66.26 ± 0.69 99.85 ± 0.05 95.40 ± 1.05 88.82 65.73 65.73
Linear probing 0.01 3,416 80.17 ± 0.85 79.73 ± 0.84 93.37 ± 0.31 81.27 ± 0.70 80.89 ± 0.39 60.23 ± 0.62 99.21 ± 0.07 91.34 ± 0.20 83.28 83.28 78.44
Prompt-tuning [27] 0.17 20,582 81.07 ± 0.59 80.02 ± 2.34 94.20 ± 0.20 81.66 ± 1.03 82.67 ± 0.34 59.62 ± 0.83 99.27 ± 0.03 90.95 ± 0.82 83.68 83.62 63.47
Attention-tuning [45] 33.8 20,792 81.99 ± 1.85 81.87 ± 1.42 94.40 ± 0.40 80.11 ± 1.92 80.67 ± 2.33 60.03 ± 1.03 99.58 ± 0.19 92.07 ± 2.48 83.68 73.74 63.34
Adapter [7] 2.03 19,360 85.84 ± 2.14 80.77 ± 7.48 95.96 ± 0.13 79.22 ± 2.26 80.76 ± 3.38 61.39 ± 1.40 99.17 ± 0.48 90.02 ± 0.91 84.14 83.41 64.61
BitFit [2] 0.12 20,382 84.31 ± 0.50 83.81 ± 1.11 94.84 ± 0.15 82.35 ± 0.62 84.77 ± 0.54 63.84 ± 1.80 99.81 ± 0.03 91.62 ± 0.95 85.67 85.63 65.11
LoRA [22] 0.68 20,970 85.84 ± 2.42 86.08 ± 0.95 95.02 ± 0.22 84.22 ± 1.98 82.26 ± 4.04 62.26 ± 1.66 99.69 ± 0.05 92.13 ± 1.12 85.94 85.69 64.93
AdaLoRA [51] 0.52 20,960 85.21 ± 0.34 86.07 ± 0.74 95.85 ± 0.24 84.23 ± 0.65 86.91 ± 0.30 62.45 ± 1.46 99.75 ± 0.04 92.25 ± 0.70 86.59 86.40 65.43
FPT [23] 1.81 1,824 84.13 ± 1.06 84.95 ± 2.01 93.88 ± 0.60 80.68 ± 1.13 90.52 ± 0.59 65.35 ± 0.94 99.70 ± 0.30 91.97 ± 1.36 86.40 85.73 83.60
FPT+ (Ours) 1.03 736 85.53 ± 1.58 86.31 ± 0.66 94.27 ± 0.32 82.91 ± 1.25 90.64 ± 0.71 66.33 ± 0.43 99.71 ± 0.18 91.30 ± 0.94 87.12 86.73 85.94
DINO (ImageNet)
Full fine-tuning 100 23,128 87.33 ± 0.77 84.06 ± 1.29 96.44 ± 0.26 84.46 ± 1.55 91.19 ± 0.32 64.70 ± 0.32 99.92 ± 0.03 95.31 ± 1.19 87.93 65.07 65.07
Linear probing 0.01 3,416 80.77 ± 0.67 81.71 ± 0.69 93.44 ± 0.16 83.96 ± 0.93 81.06 ± 0.46 63.64 ± 1.08 99.40 ± 0.01 91.84 ± 0.18 84.48 84.48 79.57
Prompt-tuning [27] 0.17 20,582 80.48 ± 0.34 80.71 ± 1.85 93.76 ± 0.11 83.41 ± 1.40 81.86 ± 0.34 63.30 ± 1.25 99.44 ± 0.14 91.97 ± 0.33 84.37 84.31 63.99
Attention-tuning [45] 33.8 20,792 82.38 ± 0.93 77.11 ± 3.29 92.60 ± 0.74 76.32 ± 4.05 75.58 ± 8.26 51.31 ± 2.63 99.57 ± 0.14 95.03 ± 0.30 81.24 71.59 61.49
Adapter [7] 2.03 19,360 87.95 ± 0.54 84.99 ± 1.89 96.02 ± 0.13 82.30 ± 2.26 77.61 ± 8.86 57.56 ± 1.48 99.79 ± 0.12 89.73 ± 4.55 84.49 83.76 64.88
BitFit [2] 0.12 20,382 86.35 ± 0.42 84.68 ± 0.42 95.11 ± 0.15 84.60 ± 1.13 84.27 ± 0.34 66.03 ± 1.53 99.82 ± 0.04 93.43 ± 0.64 86.79 86.74 65.96
LoRA [22] 0.68 20,970 85.14 ± 3.86 84.35 ± 1.27 95.97 ± 0.40 81.06 ± 6.82 $7$ 1.92 60.59 ± 2.36 99.89 ± 0.03 95.02 ± 0.70 84.97 84.72 64.20
AdaLoRA [51] 0.52 20,960 86.39 ± 0.37 84.27 ± 0.66 95.79 ± 0.25 83.74 ± 0.88 88.16 ± 0.35 64.39 ± 0.14 99.64 ± 0.08 93.29 ± 0.79 86.96 86.76 65.71
FPT [23] 1.81 1, 824 83.79 ± 0.76 84.69 ± 1.06 94.27 ± 0.18 83.05 ± 0.85 90.08 ± 0.68 66.23 ± 0.46 99.72 ± 0.05 91.55 ± 0.81 86.67 86.00 83.86
FPT+ (Ours) 1.03 736 86.52 ± 0.42 86.13 ± 0.85 93.82 ± 0.40 84.14 ± 1.65 90.50 ± 0.37 65.75 ± 0.44 99.73 ± 0.14 91.52 ± 0.72 87.26 86.87 86.08
@@ IV-B1 Experiment Setup
In this study, we employ two popular variants of ViT [15], specifically ViT-B(ase) and ViT-L(arge), as our large-scale pre-trained models, with 86 million and 304 million parameters, respectively. To evaluate the generalizability of different PETL methods, we experiment with two types of pre-trained weights: one obtained through supervised learning and the other via self-supervised learning using DINO [5]. These models, which are pre-trained on the ImageNet-21K dataset [40], are widely recognized for their robust generalizability in natural image tasks. The patch size of ViTs used in this work is 16. All methods are trained for 20 epochs with a mini-batch size of 16 on an NVIDIA RTX A6000 GPU (48GB) using PyTorch of version 2.0.0 [36]. We adopt a cosine annealing schedule for the learning rate during training. ViTs are trained using the AdamW optimizer with cross-entropy as the loss function for all datasets. Each experiment is repeated 5 times, and we report the mean ± standard deviation values.
If not specified, all compared methods run at a resolution of $512\times 512$. In the context of FPT+, the high resolution for the LPM remains $512\times 512$, while the low resolution for the side network is set to $128\times 128$. The number of fine-grained prompts is set to 16, and the reduction factor $k$ is set to 8. For important tokens selection, the top 20% of the important tokens are retained for utilization within the fusion modules. The number of layers in the side network $L_{S}$ is set to 6, which is half the number of layers in the LPM.
@@ IV-B2 Evaluation Metrics
We use the Area Under the Receiver Operating Characteristic Curve (AUC) to evaluate the classification performance on each dataset. To evaluate the performance and efficiency balance, we adopt the metrics proposed in our previous work [23], which extend from the performance-efficiency metric [32, 20]. Specifically, two metrics, namely performance-parameter-efficiency (PPE) and performance-memory-efficiency (PME), are utilized to assess the impact of the number of learnable parameters and GPU memory consumption on efficiency during transfer learning, respectively. They are defined as
@@
LoRA [22] 107.89 1014.68 20.62 295.85 288.47
AdaLoRA [51] 107.28 1027.17 20.44 292.00 280.56
Preloading 87.84 31.76 32.81 221.812 217.74
FPT 0.28 154.23 171.83 166.54 183.10
@@
TABLE III: Comparison results with state-of-the-art PETL methods using a ViT-L backbone with supervised pre-trained weights, evaluated at different input resolutions. The ’/’ symbol indicates out-of-memory conditions on a 48GB GPU. The average AUC is computed across all eight classification tasks.

### v3

@@ Abstract
The success of large-scale pre-trained models has established fine-tuning as a standard method for achieving significant improvements in downstream tasks. However, fine-tuning the entire parameter set of a pre-trained model is costly. Parameter-efficient transfer learning (PETL) has recently emerged as a cost-effective alternative for adapting pre-trained models to downstream tasks. Despite its advantages, the increasing model size and input resolution present challenges for PETL, as the training memory consumption is not reduced as effectively as the parameter usage. In this paper, we introduce Fine-grained Prompt Tuning plus (FPT+), a PETL method designed for high-resolution medical image classification, which significantly reduces the training memory consumption compared to other PETL methods. FPT+ performs transfer learning by training a lightweight side network and accessing pre-trained knowledge from a large pre-trained model (LPM) through fine-grained prompts and fusion modules. Specifically, we freeze the LPM of interest and construct a learnable lightweight side network. The frozen LPM processes high-resolution images to extract fine-grained features, while the side network employs corresponding down-sampled low-resolution images to minimize the memory usage. To enable the side network to leverage pre-trained knowledge, we propose fine-grained prompts and fusion modules, which collaborate to summarize information through the LPM’s intermediate activations. We evaluate FPT+ on eight medical image datasets of varying sizes, modalities, and complexities. Experimental results demonstrate that FPT+ outperforms other PETL methods, using only 1.03% of the learnable parameters and 3.18% of the memory required for fine-tuning an entire ViT-B model. Our code is available on https://github.com/YijinHuang/FPT.
@@
Parameter-efficient transfer learning, Memory-efficient transfer learning, Large-scale pre-trained models, High-resolution medical image classification.
@@ I Introduction
In natural images, the objects of interest typically occupy a significant portion of the image and exhibit clear and noticeable attributes (e.g., shape and color). In contrast, the diagnostic cues within medical images are often tiny and spread throughout the entire visual field, making fine-grained information crucial for accurate medical image analysis [25]. Providing such fine-grained information often requires high-resolution input images [24]. However, as illustrated in Fig. 2, this preference for high resolution comes at the cost of increased GPU memory consumption and higher training expenses. For example, a ViT with a patch size of 16 divides a 224 $\times$ 224 image into 196 patches, whereas a 512 $\times$ 512 image generates 1,024 patches. This sharp rise in computational demand makes fine-tuning LPMs both expensive and, in some cases, infeasible on devices with limited GPU memory.
@@
Additionally, considering that long input sequences from high-resolution images result in substantial intermediate activations, which can be computationally expensive, FPT+ employs important tokens selection to mitigate this issue. Based on the prior knowledge that medical images of the same modality often share similar anatomical characteristics and have small objects of interest (e.g. lesions and vessels in fundus images, and lungs in chest X-rays), FPT+ selectively fuses features from the intermediate activations to reduce GPU memory consumption. Preloading techniques are also proposed to improve the training efficiency by pre-storing the LPM’s intermediate features.
@@
  • We evaluate different LPMs for transfer learning, including ImageNet-based supervised and self-supervised LPMs, as well as a medical-specific pre-trained LPM, emphasizing the generalizability of FPT+ across different LPM scales and different pre-trained weights.
This paper extends our previous work, FPT [23]. We systematically investigate the impact of several key components in FPT and refine the entire framework to make it simpler and more efficient. Specifically, our extension is presented in several aspects: (1) We reduce the number of layers in the side network and its corresponding fusion modules to significantly decrease memory consumption; (2) We redesign the fusion module by replacing in-projector with layer normalization and align the initial hidden dimension of the fine-grained prompts with the frozen LPM; (3) All fusion modules share the same weights for the out-projector, further reducing the number of learnable parameters; (4) The new design improves the effectiveness of learning pre-trained knowledge, allowing smaller low-resolution inputs for the side network, thus further reducing memory consumption; (5) To verify the generalizability of FPT+, we validate the framework using four additional datasets; (6) We evaluate the scalability and feasibility of FPT+ across different input resolutions, pre-trained models of different scales, configurations, and weights; (7) Additional efficiency metrics, including training time, throughput, and power usage, are evaluated to comprehensively assess FPT+’s resource consumption and training efficiency; (8) Extensive ablation studies are conducted to qualitatively and quantitatively analyze the impact of each component in FPT+ on performance and efficiency.
@@ II-A Parameter-efficient Transfer Learning
Unlike other PETL methods, our proposed FPT+ aim to improve both parameter and memory efficiency during transfer learning, which is a specific and common requirement in the field of medical image analysis.
@@ II-B PETL for Medical Image Analysis
Different from our previous work, FPT [23], we simplify and enhance its framework for greater efficiency. Key improvements include reducing memory consumption through optimized side networks, fusion modules and weights sharing strategy, and enabling effective learning with smaller inputs. Unlike other PETL methods for medical image analysis, our proposed FPT+ framework aims to simultaneously address the challenge of unaffordable parameter as well as memory utilization when transferring LPMs to high-resolution medical image classification tasks.
@@ III-C Asymmetric Input
Fine-grained information holds significant importance in medical image analysis. To extract such detailed information, one straightforward approach involves increasing the resolution of the input image. However, due to the mechanics of input patchification and the self-attention layer in the ViT architecture, increasing the input resolution leads to a dramatic rise in memory consumption and training cost. For instance, in a ViT with a patch size of 16, increasing the input resolution from 224 to 512 results in an expansion of the input sequence length from 196 to 1,024. Furthermore, the dimensions of the self-attention maps in the transformer block escalate from $196^{2}$ to $1,024^{2}$. This considerable training overhead presents challenges for fine-tuning on high-resolution datasets.
Although high-resolution inputs make the training cost unfeasible, employing a pre-trained model for inference with high-resolution images remains practical. We propose here an asymmetric input strategy within the FPT+ framework. Specifically, the input images $I_{h}$ for the frozen pre-trained model are of high resolution, whereas the images fed to the learnable side network are low-resolution images $I_{l}$ that get down-sampled from the high-resolution ones. Therefore, the Equation 5 becomes
@@ III-D Fine-grained Prompts and Fusion Modules
where notations for the input source and the layer index are omitted for simplification. The $\text{CA}(\cdot)$ and $\text{LN}(\cdot)$ denote the cross-attention module and the layer normalization operation, respectively. The [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the concatenation operation, and the out-projector $f_{\text{out}}$ is a linear layer to align the hidden dimension of the fine-grained prompts to that of the side network. The out-projector is shared across all fusion modules to further increase efficiency. Notably, the number of the fine-grained prompts is considerably smaller than that of the input sequence, leading to reduced training cost. They are removed after each layer’s forward processing in the side network to maintain consistency of the sequence length across layers. The pseudo-code in Algorithm 1 for the fusion module is provided to further clarify its implementation details.
Algorithm 1 Fusion Module with Fine-grained Prompts
1:Input: Side network tokens $z_{S}$, fine-grained prompts $p$, LPM-projected keys $K_{M}$ and values $V_{M}$
2:Output: Fused token sequence $\tilde{z}_{S}$
3:$p_{\text{ln}}\leftarrow\mathrm{LN}(p)$triangleright Layer normalization
4:$p_{\text{attn}}\leftarrow\mathrm{Attn}(p_{\text{ln}},K_{M},V_{M})$triangleright Cross-attention(Q, K, V)
5:$p_{\text{out}}\leftarrow f_{\text{out}}(p_{\text{attn}})$triangleright Linear projection of attention output
6:$\tilde{p}\leftarrow p_{\text{out}}+p$triangleright Residual connection
7:$\tilde{z}_{S}\leftarrow[z_{S},\tilde{p}]$triangleright Concat prompts to side network tokens
8:return $\tilde{z}_{S}$
@@ IV-B1 Experiment Setup
In this study, we employ two popular variants of ViT [15], specifically ViT-B(ase) and ViT-L(arge), as our large-scale pre-trained models, with 86 million and 304 million parameters, respectively. To evaluate the generalizability of different PETL methods, we experiment with three types of pre-trained weights: one obtained through supervised learning on ImageNet-21K [40], another via self-supervised learning using DINO [5] pre-trained on ImageNet-21K, and the third using BioMedCLIP [52], a vision-language model pre-trained on various medical imaging modalities. These models are widely recognized for their robust generalizability in both natural and medical image tasks. The default patch size used in this work for ViTs is 16, with additional evaluations conducted on ViT-B using a patch size of 8. All methods are trained for 20 epochs with a mini-batch size of 16 on an NVIDIA RTX A6000 GPU (48GB) using PyTorch of version 2.0.0 [36]. We adopt a cosine annealing schedule for the learning rate during training. ViTs are trained using the AdamW optimizer with cross-entropy as the loss function for all datasets. Each experiment is repeated 5 times, and we report the mean $\pm$ standard deviation values.
If not specified, all compared methods run at a resolution of $512\times 512$ with ImageNet-based supervised LPM. In the context of FPT+, the high resolution for the LPM remains $512\times 512$, while the low resolution for the side network is set to $128\times 128$. The number of fine-grained prompts is set to 16, and the reduction factor $k$ is set to 8. For important tokens selection, the top 20% of the important tokens are retained for utilization within the fusion modules. The number of layers in the side network $L_{S}$ is set to 6, which is half the number of layers in the LPM.
@@ IV-B2 Evaluation Metrics
We use the Area Under the Receiver Operating Characteristic Curve (AUC) to evaluate the classification performance on each dataset. To evaluate the performance and efficiency balance, we adopt the metrics proposed in FPT[23], which extend from the performance-efficiency metric [32, 20]. Specifically, two metrics, namely performance-parameter-efficiency (PPE) and performance-memory-efficiency (PME), are utilized to assess the impact of the number of learnable parameters and GPU memory consumption on efficiency during transfer learning, respectively. They are defined as
@@
Supervised (ImageNet)
Full fine-tuning 100 23,128 89.79 $\pm$ 0.76 86.87 $\pm$ 0.53 96.65 $\pm$ 0.29 83.22 $\pm$ 1.02 92.49 $\pm$ 0.34 66.26 $\pm$ 0.69 99.85 $\pm$ 0.05 95.40 $\pm$ 1.05 88.82 65.73 65.73
Linear probing 0.01 3,416 80.17 $\pm$ 0.85 79.73 $\pm$ 0.84 93.37 $\pm$ 0.31 81.27 $\pm$ 0.70 80.89 $\pm$ 0.39 60.23 $\pm$ 0.62 99.21 $\pm$ 0.07 91.34 $\pm$ 0.20 83.28 83.28 78.44
Prompt-tuning [27] 0.17 20,582 81.07 $\pm$ 0.59 80.02 $\pm$ 2.34 94.20 $\pm$ 0.20 81.66 $\pm$ 1.03 82.67 $\pm$ 0.34 59.62 $\pm$ 0.83 99.27 $\pm$ 0.03 90.95 $\pm$ 0.82 83.68 83.62 63.47
Attention-tuning [45] 33.8 20,792 81.99 $\pm$ 1.85 81.87 $\pm$ 1.42 94.40 $\pm$ 0.40 80.11 $\pm$ 1.92 80.67 $\pm$ 2.33 60.03 $\pm$ 1.03 99.58 $\pm$ 0.19 92.07 $\pm$ 2.48 83.68 73.74 63.34
Adapter [7] 2.03 19,360 85.84 $\pm$ 2.14 80.77 $\pm$ 7.48 95.96 $\pm$ 0.13 79.22 $\pm$ 2.26 80.76 $\pm$ 3.38 61.39 $\pm$ 1.40 99.17 $\pm$ 0.48 90.02 $\pm$ 0.91 84.14 83.41 64.61
BitFit [2] 0.12 20,382 84.31 $\pm$ 0.50 83.81 $\pm$ 1.11 94.84 $\pm$ 0.15 82.35 $\pm$ 0.62 84.77 $\pm$ 0.54 63.84 $\pm$ 1.80 99.81 $\pm$ 0.03 91.62 $\pm$ 0.95 85.67 85.63 65.11
LoRA [22] 0.68 20,970 85.84 $\pm$ 2.42 86.08 $\pm$ 0.95 95.02 $\pm$ 0.22 84.22 $\pm$ 1.98 82.26 $\pm$ 4.04 62.26 $\pm$ 1.66 99.69 $\pm$ 0.05 92.13 $\pm$ 1.12 85.94 85.69 64.93
AdaLoRA [51] 0.52 20,960 85.21 $\pm$ 0.34 86.07 $\pm$ 0.74 95.85 $\pm$ 0.24 84.23 $\pm$ 0.65 86.91 $\pm$ 0.30 62.45 $\pm$ 1.46 99.75 $\pm$ 0.04 92.25 $\pm$ 0.70 86.59 86.40 65.43
FPT [23] 1.81 1,824 84.13 $\pm$ 1.06 84.95 $\pm$ 2.01 93.88 $\pm$ 0.60 80.68 $\pm$ 1.13 90.52 $\pm$ 0.59 65.35 $\pm$ 0.94 99.70 $\pm$ 0.30 91.97 $\pm$ 1.36 86.40 85.73 83.60
FPT+ (Ours) 1.03 736 85.53 $\pm$ 1.58 86.31 $\pm$ 0.66 94.27 $\pm$ 0.32 82.91 $\pm$ 1.25 90.64 $\pm$ 0.71 66.33 $\pm$ 0.43 99.71 $\pm$ 0.18 91.30 $\pm$ 0.94 87.12 86.73 85.94
DINO (ImageNet)
Full fine-tuning 100 23,128 87.33 $\pm$ 0.77 84.06 $\pm$ 1.29 96.44 $\pm$ 0.26 84.46 $\pm$ 1.55 91.19 $\pm$ 0.32 64.70 $\pm$ 0.32 99.92 $\pm$ 0.03 95.31 $\pm$ 1.19 87.93 65.07 65.07
Linear probing 0.01 3,416 80.77 $\pm$ 0.67 81.71 $\pm$ 0.69 93.44 $\pm$ 0.16 83.96 $\pm$ 0.93 81.06 $\pm$ 0.46 63.64 $\pm$ 1.08 99.40 $\pm$ 0.01 91.84 $\pm$ 0.18 84.48 84.48 79.57
Prompt-tuning [27] 0.17 20,582 80.48 $\pm$ 0.34 80.71 $\pm$ 1.85 93.76 $\pm$ 0.11 83.41 $\pm$ 1.40 81.86 $\pm$ 0.34 63.30 $\pm$ 1.25 99.44 $\pm$ 0.14 91.97 $\pm$ 0.33 84.37 84.31 63.99
Attention-tuning [45] 33.8 20,792 82.38 $\pm$ 0.93 77.11 $\pm$ 3.29 92.60 $\pm$ 0.74 76.32 $\pm$ 4.05 75.58 $\pm$ 8.26 51.31 $\pm$ 2.63 99.57 $\pm$ 0.14 95.03 $\pm$ 0.30 81.24 71.59 61.49
Adapter [7] 2.03 19,360 87.95 $\pm$ 0.54 84.99 $\pm$ 1.89 96.02 $\pm$ 0.13 82.30 $\pm$ 2.26 77.61 $\pm$ 8.86 57.56 $\pm$ 1.48 99.79 $\pm$ 0.12 89.73 $\pm$ 4.55 84.49 83.76 64.88
BitFit [2] 0.12 20,382 86.35 $\pm$ 0.42 84.68 $\pm$ 0.42 95.11 $\pm$ 0.15 84.60 $\pm$ 1.13 84.27 $\pm$ 0.34 66.03 $\pm$ 1.53 99.82 $\pm$ 0.04 93.43 $\pm$ 0.64 86.79 86.74 65.96
LoRA [22] 0.68 20,970 85.14 $\pm$ 3.86 84.35 $\pm$ 1.27 95.97 $\pm$ 0.40 81.06 $\pm$ 6.82 $7$$\pm$ 1.92 60.59 $\pm$ 2.36 99.89 $\pm$ 0.03 95.02 $\pm$ 0.70 84.97 84.72 64.20
AdaLoRA [51] 0.52 20,960 86.39 $\pm$ 0.37 84.27 $\pm$ 0.66 95.79 $\pm$ 0.25 83.74 $\pm$ 0.88 88.16 $\pm$ 0.35 64.39 $\pm$ 0.14 99.64 $\pm$ 0.08 93.29 $\pm$ 0.79 86.96 86.76 65.71
FPT [23] 1.81 1,824 83.79 $\pm$ 0.76 84.69 $\pm$ 1.06 94.27 $\pm$ 0.18 83.05 $\pm$ 0.85 90.08 $\pm$ 0.68 66.23 $\pm$ 0.46 99.72 $\pm$ 0.05 91.55 $\pm$ 0.81 86.67 86.00 83.86
FPT+ (Ours) 1.03 736 86.52 $\pm$ 0.42 86.13 $\pm$ 0.85 93.82 $\pm$ 0.40 84.14 $\pm$ 1.65 90.50 $\pm$ 0.37 65.75 $\pm$ 0.44 99.73 $\pm$ 0.14 91.52 $\pm$ 0.72 87.26 86.87 86.08
BioMedCLIP
Full fine-tuning 100 23,128 84.40 $\pm$ 0.60 83.23 $\pm$ 0.94 96.52 $\pm$ 0.02 81.61 $\pm$ 0.80 88.05 $\pm$ 0.44 63.56 $\pm$ 2.46 99.80 $\pm$ 0.04 96.87 $\pm$ 0.32 86.52 64.03 64.03
Linear probing 0.01 3,416 78.98 $\pm$ 0.28 77.38 $\pm$ 0.78 93.65 $\pm$ 0.04 75.68 $\pm$ 1.76 78.42 $\pm$ 0.13 62.89 $\pm$ 0.29 99.50 $\pm$ 0.05 92.13 $\pm$ 0.51 82.23 82.23 77.45
Prompt-tuning [27] 0.17 20,582 79.36 $\pm$ 0.36 77.22 $\pm$ 0.91 93.84 $\pm$ 0.10 76.49 $\pm$ 0.69 78.94 $\pm$ 0.12 62.24 $\pm$ 0.79 99.56 $\pm$ 0.03 92.05 $\pm$ 0.45 82.39 82.33 62.49
Attention-tuning [45] 33.8 20,792 80.46 $\pm$ 1.67 77.58 $\pm$ 2.50 94.28 $\pm$ 0.65 77.79 $\pm$ 3.24 79.09 $\pm$ 0.36 60.19 $\pm$ 4.38 99.72 $\pm$ 0.06 92.12 $\pm$ 2.08 83.24 73.35 63.00
Adapter [7] 2.03 19,360 84.16 $\pm$ 0.62 78.15 $\pm$ 0.60 95.69 $\pm$ 0.28 79.17 $\pm$ 2.39 81.63 $\pm$ 2.19 61.83 $\pm$ 1.54 99.56 $\pm$ 0.25 90.32 $\pm$ 4.47 83.71 82.98 64.28
BitFit [2] 0.12 20,382 81.78 $\pm$ 0.72 82.73 $\pm$ 1.14 94.92 $\pm$ 0.10 82.20 $\pm$ 1.06 80.78 $\pm$ 0.41 65.17 $\pm$ 1.29 99.61 $\pm$ 0.12 94.92 $\pm$ 0.98 84.87 84.83 64.50
LoRA [22] 0.68 20,970 81.65 $\pm$ 0.53 79.91 $\pm$ 0.46 95.94 $\pm$ 0.22 76.82 $\pm$ 2.83 81.01 $\pm$ 0.59 62.43 $\pm$ 1.63 99.78 $\pm$ 0.03 94.59 $\pm$ 1.49 83.67 83.42 63.22
AdaLoRA [51] 0.52 20,960 83.78 $\pm$ 0.54 80.73 $\pm$ 1.39 95.41 $\pm$ 0.22 80.07 $\pm$ 0.99 84.03 $\pm$ 0.25 64.01 $\pm$ 1.00 99.72 $\pm$ 0.02 94.66 $\pm$ 0.89 85.24 85.05 64.41
FPT [23] 1.81 1,824 82.48 $\pm$ 0.36 79.81 $\pm$ 2.12 94.33 $\pm$ 0.64 83.82 $\pm$ 0.72 86.91 $\pm$ 0.43 64.06 $\pm$ 0.91 99.67 $\pm$ 0.08 92.26 $\pm$ 0.76 85.30 84.64 82.53
FPT+ (Ours) 1.03 736 82.38 $\pm$ 0.25 83.73 $\pm$ 0.30 94.50 $\pm$ 0.31 83.88 $\pm$ 0.83 88.53 $\pm$ 0.97 65.98 $\pm$ 0.82 99.84 $\pm$ 0.06 94.57 $\pm$ 0.14 86.81 86.42 85.64
@@
LoRA [22] 107.89 1,014.68 20.62 295.85 288.47
AdaLoRA [51] 107.28 1,027.17 20.44 292.00 280.56
Preloading 87.84 31.76 32.81 221.81 217.74
FPT [23] 0.28 154.23 171.83 166.54 183.10
@@
TABLE III: Comparison results with state-of-the-art PETL methods with the ImageNet-based supervised LPM, evaluated at different input resolutions. The ’/’ symbol indicates out-of-memory conditions on a 48GB GPU. The average AUC is computed across all eight classification tasks.